Minutes of the Tracker Meeting

August 19, 2004

 

Action Items:

-        Ronaldo: talk to Plyform about scheduling a visit of the anomaly review team

-        Sandro: respond to Neilís question list

-        Arthur: submit a proposal for a bonding test to be done at SLAC next week

-        Sandro: review the status of the top tray and what can be done in the near term to make it usable for Tower-A assembly

-        Dave: check with Martin on the status of the alignment plan

-        Jeff: get all the hardware shipped to Pisa before the end of the week

-        Hiro: test the C0 cables with MCMs, including the bias leakage current

 

1) Bias-circuit bonding anomaly review team

-        Schedule: The ďTiger TeamĒ will be ready for work on Monday morning, August 30, in Pisa.They will be departing Friday morning of the same week. 1 day will be spent at Plyform, probably Tuesday or Wednesday (Ronaldo will talk to Plyform).Ronaldo said that the prospective Italian consultant is on holidays, so no contact had been made yet.

-        NDAs and access to the detailed Plyform procedures: Ronaldo suggests not to open this issue unless needed. Sandro thinks there is not much to know that we donít already know. Sandro will respond to Neilís question list.

-        Sandroís question: is there any quantitative method to evaluate the cleanliness of the tiles, so they could study it next week? Ben suggested simply a traditional water break test: deionized water is dropped onto the surface to see if it spreads smoothly; if it beads up like on a waxed car, then it is not a good surface for bonding.

-        Bonding test using etch (Mil recipe): Arthur will make a proposal for a test at SLAC.Robert noted that Plyform has an etch procedure that they use for Agile trays.

-        Neil asked about the typical bond thickness.Sandro said that it is close to 50 microns (2 mils).

-        Neil said that he is trying to identify a time for a telecon on the analysis. John Ku said that he will have some results summarized soon, hopefully today. Ben recommended a strain based criteria, rather than stress, because of nonlinearities of the adhesive, especially with temperature.He stressed that we donít have enough information on stress/strain curves versus temperature for this adhesive to do the analysis in this way.

 

2) Tower A tray status

-        Completion of the work at G&A: G&A has stopped work and gone on vacation.

-        Status: 16 trays in Pisa; 4 are shorted on one side; 2 have no ladders but have MCMs.

-        The 3 missing trays are set aside for investigation. 1 tray is 200 microns small in both directions.

-        We discussed the possibility of assembling working MCMs on the 3 trays. Sandro will look at the status of the top tray in particular and see what is possible.The tracker team would like to maintain the old schedule for Tower A assembly (starting August 26) with the trays as is, as long as all have MCMs mounted.SLAC and NASA management have yet to approve, however.

-        Stacked tray testing plans, schedule, procedure release status PS-03361.Luca Baldini said that it should be done already and ready for release.Robert will ask Luca on Monday.

 

3) Tower Assembly PRR readiness

-        Assembly procedure PS-01854: still in work; to be released next week

-        Alignment plan/procedure PS-03566: Dave will check the status with Martin

-        Arthur: interface assembly procedure

-        Electrical test procedure TD-00191: check with Luca

-        Cosmic-ray test procedure PS-01855: apparently this number is not being used

-        Packaging, handling, transportation PS-04525: in work, with the first draft to come out next week

-        PRR presentations: Emilio will prepare charts; Luca should also have some charts

 

4) Status of SLAC deliverables for Tower A

-        A large number of items need to be shipped from SLAC this week

-        lifting fixture: ready to ship

-        inner container: proof loading complete

-        cable holding plate: ready to ship

-        vibe fixture: ready to ship

-        screws and pins: ready to ship

-        I/F hardware: all ready, except the flight studs, which are still getting coated. Arthur will carry the studs to Pisa.

-        cable bending tool and cable installation testing: Arthur will carry a mockup to Pisa for testing the installation.

-        corner brackets and clips and nests: corner brackets are coming out of plating today

-        C0 cables: Robert said that at least 3 new ones are in the clean room, and Robert will hand carry them to Pisa.Luca wants SLAC to include connector savers with them. Robert asked Hiro to check whether he can test them with MCMs before Monday.

 

5) MCM issues

-        pitch adapter cracking: SLAC is inspecting MCMs post-burn-in to try to avoid the situation with one MCM on a tray that has 25 broken traces.So far, some MCMs have been found with many apparent cracks, but almost all of the traces still conducted.Further study is in progress.Work is still going on to make an improved pitch adapter.The first attempt failed.The second attempt will have the solder mask extend all the way around the bend region, in which case Teledyne must improve the alignment (some improvement already made) to insure that the wire bond pads donít get lost.

-        HV shorts: 6 MCMs have been found in the post-burn-in test to have HV shorts internal to the PWB between the HV and AVDDA planes.Further investigation is in progress.

-        HV capacitors: the Tracker data on MCMs do not show any problem with leakage current, except that there is nearly always some decrease in current from the burn-in.No MCMs have shorted so far due to capacitor failure.The PCB apparently approved flight use of the existing MCMs with the questionable capacitors.Novacap is producing a replacement lot that we will use at Teledyne when available.